OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24> @Stiwan

OpenStudy (anonymous):

do u knw dot product of two vectors?

OpenStudy (anonymous):

noi sorry

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

@SandeepReddy

OpenStudy (anonymous):

you use the dot product too vectors that are perpendicular i unit vector (dot) j unit vector = 0 because A dot B= |A||B|cos X where x is the angle between the vectors A dot B = Ax*Bx + Ay*By thats for two dimension vector if you have two vectors you can find the coeficients that make them perpendicular by setting the dot product to 0 ai + 3j = A 6i + 1j = B A dot B = 6 * a + 3 * 1 = 6a + 3 setting it to zero 6a + 3 = 0 implys 6a = -3 a = -1/2 so the vector A is -1/2i + 3j when A is perpendicular to B The line is parallel hmm have to think thats the CROSS PRODUCT WHEN you take the cross product of two vectors that gives A x B =|A||B|sin X the function will be zero when the angle X IS 0 OR 180 SINCE THE ANGLE x is the angle between vector A and B when the cross product is zero the lines are parallel you can calculate the cross product | Ax Ay | | Bx By | = AxBy - AyBx HOPE THAT HELPS ... SUMMARIZE DOT PRODUCT two vectors = 0 when perpendicular cross product of two vectors is 0 when parallel

OpenStudy (anonymous):

thanks so much !

OpenStudy (anonymous):

My pleasure!